# Molecular clouds: do they deserve a non-Gaussian description?

### v1

@@ Acknowledgements.
This work reused the Describable Textures Dataset [^15] https://www.robots.ox.ac.uk/~vgg/data/dtd/.
Appendix A Other datasets
A.1 Numerical simulations
Figure 18: Overview of the simulations. One snapshot is shown for each model (”Hydro high res”, ”MHD”, and ”MHD high B”), and the maps show the central $33\,\mathrm{pc}\times 33\,\mathrm{pc}$ field for gas column density integrated along the $y$ axis.
@@ A.1 Numerical simulations
A.2 logFBM models
To perform our analysis of observational data, we also consider purely synthetic, parametric models of column density maps, derived from exponentiations of fractional Brownian motions (fBm). Such fields have been previously studied by the ISM community (Elmegreen 2002; Brunt & Heyer 2002; Miville-Deschênes et al. 2007; Levrier et al. 2018). It is indeed a convenient class of models that allows to simultaneously reproduce exactly the log-normal one-point properties (encountered in quiescent regions) and power-law power spectra to a good approximation. An example is given in Fig. 1.
@@
such that $\sigma F_{\beta}\star Z+\mu$ is a real valued Gaussian process with mean $\mu$, variance $\sigma^{2}$ and $\beta$-decaying power spectrum power-law.
The parameters $(\mu,\sigma,\beta)$ are fitted from the observations. More precisely, for each $512\times 512$ column density patch that has no negative pixel (that is about $87\%$ of the 550 observational patches), we estimate the mean $\mu$, variance $\sigma^{2}$ and spectral index $\beta$ on the logarithm of the column density patch, expressed in $10^{22}\,\mathrm{cm^{-2}}$ units. This procedure leads to a distribution of $\sim 480$ estimated parameters triplets $(\mu,\sigma,\beta)$ that is reported in Fig. 19. For each triplet, we then sample one $512\times 512$ logFBM patch. Contrary to the other patches considered in this paper (observations, simulations, everyday textures), these samples have periodic boundary conditions, by construction. However, their $256\times 256$ sub-patches, that are the only ones on which the various summary statistics are applied, do not, just as the rest of the data.
We dub these synthetic maps logFBM models, to recall that the statistics of their logarithms are those of the Gaussian, fBm random fields. In the literature, emission maps are often produced by integrating a 3D generated field. In this paper, for simplicity and computational efficiency, we directly generate 2D fields. Even though integrated 3D fields could be expected to better model MCs’ column density maps, they do not differ much in morphology from 2D models and both are far from reproducing MCs’ morphology. Thus, the 2D logFBM models used here are suited to the scope of this paper.
@@
A.3 Describable Textures Dataset (DTD)
To better understand the peculiarities of MCs’ statistical properties, we compare them to a very wide set of everyday textures: the Describable Textures Dataset (DTD) (Cimpoi et al. 2014). This dataset is in open access [^17] https://www.robots.ox.ac.uk/~vgg/data/dtd/. In this paper, we keep only the images with smaller edge size larger than 512 pixels. We then crop each one of these $\sim 1050$ resulting pictures around its center into a $512\times 512$ patch. Then each patch is converted to gray-scale by summing its three RGB channels. These maps have integer values that range from 0 to $3\times 255=765$ that we finally divide by a factor $10^{3}$ to match the typical values of the other datasets.
Appendix B Apodization
The $256\times 256$ sub-patches on which we perform 2D Fourier transform do not have periodic boundary conditions (PBC). Thus, before applying such transform during power spectrum estimation on a given sub-patch $x$, we first apodize it as follows:
$x\mapsto rw\cdot(x-\langle x\rangle_{\@vec{u}}),$
where $w\cdot$ denotes the pixel-wise multiplication by the window $w$, reported in Fig. 20 and $r$ is a scalar factor depending on the input map $x$ such that the output apodized map has same variance as the input. The resulting map does not share the same mean with the input but this property is not regarded by the Fourier-based statistics used here.
Appendix C Srivastava & Du test statistic
The test statistic $d^{2}_{\phi}(x_{i},x_{j})$ we introduce in Eq. 2 is derived from Srivastava & Du (2008). We apply this test on the two collections of summary statistics $\{\phi(x_{i}^{(l)})\}_{1\leq l\leq N_{i}}$ and $\{\phi(x_{j}^{(l)})\}_{1\leq l\leq N_{j}}$ computed on the sub-patches (indexed by $l$) of the patch $x_{i}$ and the patch $x_{j}$. In our case, $N_{i}=N_{j}=4$. We recall Eq. 2 here:
@@ Appendix C Srivastava & Du test statistic
$\begin{cases}R\equiv D_{S}^{-1/2}SD_{S}^{-1/2},\\ c_{p,n}:=1+trR^{2}/p^{3/2},\\ \alpha\equiv\frac{N_{i}N_{j}}{N_{i}+N_{j}}\frac{1}{\sqrt{2(trR^{2}-p^{2}/n)c_{p,n}}},\\ \beta\equiv\frac{N_{i}+N_{j}}{N_{i}N_{j}}\frac{np}{n-2}.\end{cases}$
Appendix D Why taking the logarithm of some standard statistics?
In theory, because the mapping $\psi\mapsto\log\psi$ is invertible, the same amount of information should be contained in $\phi(x)$ and in $\log\phi(x)$ (when the latter is well defined). However, for the compatibility diagnostic we chose, these two options might perform differently. Such issues are not specific to this test and also occur with standard diagnostics such as Fisher analysis Park et al. (2023).
First, this non-linear mapping can change the fulfillment level of the test assumptions. In the Srivastava & Du test used here, the two input distributions that are to be compared are assumed to be normally distributed and have equal variance. In many cases, typically when $\phi(x)>0$, taking the logarithm of $\phi(x)$ decreases the deviation to normality and decreases the relative deviations between the variances of $\phi(x_{i})$ and $\phi(x_{j})$.
@@ Appendix D Why taking the logarithm of some standard statistics?
Figure 21: Effect of the log rescaling after computing variance statistics on observation data. Provided that the assumptions of the test are not too discarded in both cases for these results to be appropriately interpreted on both axes, this shows that $\big{\langle}\big{[}x-\langle x\rangle_{\@vec{u}}\big{]}^{2}\big{\rangle}_{\@vec{u}}$ is strictly more degenerate than $\log\,\big{\langle}\big{[}x-\langle x\rangle_{\@vec{u}}\big{]}^{2}\big{\rangle}_{\@vec{u}}$.

### v2

@@ Acknowledgements.
We thank the anonymous referee for his careful reading of the manuscript, detailed comments and useful recommendations that helped us to improve the quality of the paper.
@@
This work used the Describable Textures Dataset [^16] https://www.robots.ox.ac.uk/~vgg/data/dtd/.
Appendix A Further details on the datasets
A.1 Observations
Fig. 14 presents the footprints on the sky of the set of observations used in this work.
Figure 14: Footprints of the Herschel Gould Belt Survey (HGBS) fields used in this study, overlaid on the total thermal dust intensity at $353\,\mathrm{GHz}$ from the GNILC (Remazeilles et al. 2011) variable resolution data of Planck (Planck Collaboration 2020).
A.2 Numerical simulations
Figure 15: Overview of the simulations. One snapshot is shown for each model (”Hydro high res”, ”MHD”, and ”MHD high B”), and the maps show the central $33\,\mathrm{pc}\times 33\,\mathrm{pc}$ field for $\mathrm{H_{2}}$ column density integrated along the $y$ axis.
@@ A.2 Numerical simulations
A.3 logFBM models
To perform our analysis of observational data, we also considered purely synthetic, parametric models of column density maps, derived from exponentiations of FBMs. Such fields have been previously studied by the ISM community (Elmegreen 2002; Brunt & Heyer 2002; Miville-Deschênes et al. 2007; Levrier et al. 2018). It is indeed a convenient class of models that allows us to simultaneously reproduce exactly the log-normal one-point properties (encountered in quiescent regions) and power-law power spectra to a good approximation. An example is given in Fig. 1.
@@
such that $\sigma F_{\beta}\star Z+\mu$ is a real valued Gaussian process with mean $\mu$, variance $\sigma^{2}$ and $\beta$-decaying PS power-law.
The parameters $(\mu,\sigma,\beta)$ are fitted from the observations. More precisely, for each $512\times 512$ column density patch that has no negative pixel (that is about $87\%$ of the 550 observational patches), we estimate the mean $\mu$, variance $\sigma^{2}$ and spectral index $\beta$ on the logarithm of the column density patch, expressed in $10^{22}\,\mathrm{cm^{-2}}$ units. This procedure leads to a distribution of $\sim 480$ estimated parameters triplets $(\mu,\sigma,\beta)$ that is reported in Fig. 16. For each triplet, we then sample one $512\times 512$ logFBM patch. Contrary to the other patches considered in this paper (observations, simulations, everyday textures), these samples have PBCs, by construction. However, their $256\times 256$ subpatches, that are the only ones on which the various summary statistics are applied, do not, just as the rest of the data.
We dub these synthetic maps logFBM models, to recall that the statistics of their logarithms are those of FBM fields. In the literature, emission maps are often produced by integrating a 3D generated field. In this paper, for simplicity and computational efficiency, we directly generate 2D fields. Even though integrated 3D fields could be expected to better model MCs’ column density maps, they do not differ much in morphology from 2D models and both are far from reproducing MCs’ morphology. Thus, the 2D logFBM models used here are suited to the scope of this paper.
@@
A.4 Describable textures dataset
To better understand the peculiarities of MCs’ statistical properties, we compare them with a very wide set of everyday textures: the DTD (Cimpoi et al. 2014). This dataset is openly accessible [^18] https://www.robots.ox.ac.uk/~vgg/data/dtd/. In this paper, we keep only images with a smaller edge size larger than 512 pixels. We then crop these $\sim 1050$ resulting pictures into $512\times 512$ patches. Next, each patch is converted to grayscale by summing its three RGB channels. These maps have integer values ranging from 0 to $3\times 255=765$ that we finally divide by a factor $10^{3}$ to match the typical values of the other datasets.
Appendix B Srivastava and Du test statistic
The test statistic $d^{2}_{\phi}(x_{i},x_{j})$ we introduce in Eq. 2 is derived from Srivastava & Du (2008). We apply this test on the two collections of summary statistics $\{\phi(x_{i}^{(l)})\}_{1\leq l\leq N_{i}}$ and $\{\phi(x_{j}^{(l)})\}_{1\leq l\leq N_{j}}$ computed on the subpatches (indexed by $l$) of the patch $x_{i}$ and the patch $x_{j}$. In our case, $N_{i}=N_{j}=4$. We recall Eq. 2 here:
@@ Appendix B Srivastava and Du test statistic
$\begin{cases}R\equiv D_{S}^{-1/2}SD_{S}^{-1/2},\\ c_{p,n}\equiv 1+\operatorname{tr}R^{2}/p^{3/2},\\ \alpha\equiv\frac{N_{i}N_{j}}{N_{i}+N_{j}}\frac{1}{\sqrt{2(\operatorname{tr}R^{2}-p^{2}/n)c_{p,n}}},\\ \beta\equiv\frac{N_{i}+N_{j}}{N_{i}N_{j}}\frac{np}{n-2}.\end{cases}$
Appendix C Rationale for using the logarithm of some standard statistics
In theory, because the mapping $\psi\mapsto\log\psi$ is invertible, the same amount of information should be contained in $\phi(x)$ and in $\log\phi(x)$ (when the latter is well defined). However, for the compatibility diagnostic we chose, these two options might perform differently. Such issues are not specific to this test and also occur with standard diagnostics such as Fisher analysis (Park et al. 2023).
First, this nonlinear mapping can change the fulfillment level of the test assumptions. In the Srivastava and Du test used here, the two input distributions to be compared are assumed to be normally distributed and with equal variance. In many cases, typically when $\phi(x)>0$, taking the logarithm of $\phi(x)$ decreases the deviation to normality and decreases the relative deviations between the variances of $\phi(x_{i})$ and $\phi(x_{j})$.
@@ Appendix C Rationale for using the logarithm of some standard statistics
Figure 17: Effect of logarithmic rescaling after computing variance statistics on observation data. Provided that the assumptions of the test are not too discarded in both cases for these results to be appropriately interpreted on both axes, this shows that $\big{\langle}\big{[}x-\langle x\rangle_{\@vec{u}}\big{]}^{2}\big{\rangle}_{\@vec{u}}$ is strictly more degenerate than $\log\,\big{\langle}\big{[}x-\langle x\rangle_{\@vec{u}}\big{]}^{2}\big{\rangle}_{\@vec{u}}$.
Appendix D Apodization
The $256\times 256$ subpatches on which we perform 2D Fourier transform do not have PBCs. Thus, before applying such transform during PS estimation on a given subpatch $x$, we first apodize it as follows:
$x\mapsto rw\cdot(x-\langle x\rangle_{\@vec{u}}),$
where $w\cdot$ denotes the pixel-wise multiplication by the window $w$, reported in Fig. 18 and $r$ is a scalar factor depending on the input map $x$ such that the output apodized map has same variance as the input.
The resulting map does not share the same mean with the input but this property is not regarded by the Fourier-based statistics used here.
Appendix E Additional examples of degeneracies
We show six examples of pairs of patches in Fig. 19 that have compatible log-Gaussian statistics but incompatible RWST statistics, as shown in Fig. 20.
Figure 19: Examples of log-Gaussian degeneracies. Six pairs of $512\times 512$ patches are chosen, whose locations on the scatter plot of Fig. 10.b are given there by the red stars. The column density maps are shown in units of $10^{20}\text{cm}^{-2}$. For each patch, we report: $s$ the subsampling factor from the original 3″/pix map, $d$ and $b$ the approximated distance and Galactic latitude of the cloud. If a pair has patches $(i,j)$ with incompatible pixel sizes according to the following criterion $\max\{\frac{s_{i}d_{i}}{s_{j}d_{j}},\frac{s_{j}d_{j}}{s_{i}d_{i}}\}\geq 3/2$, we color its labels in red.
Figure 20: Statistics for the examples of log-Gaussian confusions shown in Fig. 19. In each row, the orange filled line (resp. band) corresponds to the mean (resp. standard deviation) of the statistics computed over the four $256\times 256$ subpatches of the top patch of each pair of Fig. 19, and the corresponding blue lines and areas refer to the bottom patch of the pair. The top row corresponds to the RWST, starting with $S_{1}^{Iso}[j]$ coefficients with four increasing scales $j$, then $\langle S_{2}^{Iso1}\rangle_{j_{2}-j_{1}}[\delta]$ coefficients with three increasing scale ratios $\delta$ and finally $\langle S_{2}^{Iso2}\rangle_{j_{2}-j_{1}}[\delta]$ coefficients with the same three scale ratios. The bottom row represents in the following order: mean of log followed by PS of log with six decreasing scales. The second and third rows show the offsets of these statistics with respect to the mean of the two.
Appendix F Lower bound of the gap between observations and simulations
In Sect. 6.3, we investigate whether observations and simulations do intersect. Ideally, we would like to check whether the set of observations’ moments $\{\bar{\mu}_{i}\}_{i\in\text{obs}}$ overlaps that of simulations $\{\bar{\mu}_{j}\}_{j\in\text{sim}}$, where $\bar{\mu}_{i}\equiv\mathbb{E}[\phi_{\text{final}}(x_{i})]$ stands for the expected value over a given process $i$. In practice, we would aim at retrieving the minimal distance
between $\bar{\mu}_{i}$ and $\bar{\mu}_{j}$ over (OBS, SIM) pairs $(i,j)$.
However, the distance $d^{2}_{\text{obs}}$ introduced previously is a statistical estimator so that, in general,
$d^{2}_{ij}\neq\bar{d}^{2}_{ij},$
and in particular: $\mathbb{E}[d^{2}_{ij}]>\bar{d}^{2}_{ij}$. Indeed, the variance of the estimator $\hat{\mu}_{i}$ biases $d^{2}_{ij}$ with respect to $\bar{d}^{2}_{ij}$:
$\mathbb{E}[d^{2}_{ij}]\equiv\bar{d}^{2}_{ij}+b_{ij}.$ (23)
This non-negative bias, which boils down to
prevents us from interpreting directly the value of $d^{2}_{ij}$ as $\bar{d}^{2}_{ij}$. Furthermore, its value increases with the amplitude of the fluctuations of $\hat{\mu}_{i}-\hat{\mu}_{j}$ and can thus change depending on the pair $(i,j)$ considered. This dependency on the pair also prevents us, without further check, from actually comparing $\bar{d}^{2}$ between different pairs based on the estimations $d^{2}$.
  • Find an observation $i$ and a simulation $j$ that minimize the biased estimation $d^{2}_{ij}$. Such pairs are already reported in Fig. 12 and are good candidates to minimize $\bar{d}^{2}_{ij}$. The goal then becomes to compare $\bar{d}^{2}_{ij}$ with the minimal value $\bar{d}^{2}_{ii^{\prime}}$, for $i^{\prime}$ another observation, independent of $i$. But $\bar{d}^{2}$ is unknown. Instead:
  • If, in addition the observation patch $x_{i^{\prime}}$ is such that spatial fluctuations of $\phi_{\text{final}}$ over its subpatches are at least of the order of those estimated from the subpatches of the simulation patch $x_{j}$, then based on Eq. 24, we have $b_{ii^{\prime}}\gtrapprox b_{ij}$. This allows us then to use $d^{2}_{ij}-d^{2}_{ii^{\prime}}$ as an estimated lower bound for $\bar{d}^{2}_{ij}-\bar{d}^{2}_{ii^{\prime}}$.
We report in Fig. 13 such independent [^19] Indeed, only one of these pairs is made of patches retrieved from the same MC (Ophiuchus), but these patches do not overlap. pairs $(i,i^{\prime})$. In the first row, we fix $i$ associated with the patch from Ophiuchus $x_{i}$ that is the closest observation to simulations ($d^{2}_{ij}=7$, Fig. 12). For this choice of patch, the reported distances $d^{2}_{ii^{\prime}}$ of neighboring observations $i^{\prime}$ are rather similar to $d^{2}_{ij}$. However, in the second row, where $i$ is associated with the patch from Aquila that is the second closest observation to simulations (as shown in Fig. 12), we see other observations $i^{\prime}$, such as Serpens or Orion B, that are closer than the closest simulation: $d^{2}_{ii^{\prime}}=3<d^{2}_{ij}=8$. We checked that the fluctuations of $i^{\prime}$ are comparable to those of $j$ so that $b_{ii^{\prime}}\gtrapprox b_{ij}$. This implies then $\bar{d}^{2}_{ij}-\bar{d}^{2}_{ii^{\prime}}\gtrapprox 5$ while $\bar{d}^{2}_{ii^{\prime}}\leq d^{2}_{ii^{\prime}}=3$. Hence, $\bar{d}^{2}_{ij}/\bar{d}^{2}_{ii^{\prime}}\gtrapprox 8/3$. This second case shows that one of the closest (OBS, SIM) pairs is definitely further apart than this specific observation is with other observations, at least according to $\phi_{\text{final}}$. This example also supports that the potential bias on the $d^{2}_{\text{obs}}$ estimators remains moderate for the study of (OBS, SIM) pairs ($\bar{d}^{2}_{ij}\gtrapprox 5$ and $d^{2}_{ij}=8$ implies $b_{ij}\lessapprox 3$).